OpenStudy (anonymous):

olve the equation. Enter your answer like this example. j=3 j + 4 = - 17

jimthompson5910 (jim_thompson5910):

What do you think you have to do here?

OpenStudy (anonymous):

To solve for j we want it to be by itself. Let's keep it on the left side so now we have to get rid of the +4. In order to do that we have to subtract both sides by 4 \[j + 4 = - 17 \] \[j + 4-4 = - 17-4 \] What is j?